n the spectrum, the relative heights of the integrals are 8.5cm, 8.5cm and 13.5cm( from left to right). what is the inference about the number of protons under each signal
Answer #1

Answer :

Given 3 integrals minimum ratio min. no. of protons under each signal
8.5 8.5/8.5 = 1 2
8.5 8.5/8.5= 1 2
13.5 13.5/8.5 = 1.58 3

Since minimum ratio for 13.5 cm integral is 1.58 . Since no. of protons can't be fractional. So, the minimum ratio must be multiplied by the 2.